OpenStudy (anonymous):

4.03 Algebra 1 Math Help!

OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1.Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. 2.A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. 3.Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. 4.In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents

OpenStudy (anonymous):

help @spac3cadet ?

OpenStudy (anonymous):

f(x)=a.ekt with k is growth rate each year. t is time (measured in year) Principal population for 15 years can be calculated by substituting t = 15 into the functional model. Rate of growth k=1tlnf(x)a If we know principal population at 15 years after, then we can also know the growth rate k

OpenStudy (anonymous):

g(x) = 200(1.03)^(2x) g(x) = 200( 1.03^2 )^x ... using the rule: a^(b*c) = (a^b)^c g(x) = 200( 1.0609 )^x

OpenStudy (anonymous):

idk.

OpenStudy (anonymous):

wait so what is the equation for #1?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i am sure i walked through this exact question before must be a virus we can do this \[g(x) = 200(1.03)^{2x}\]. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent in one second

OpenStudy (anonymous):

\[(1.03)^{2x}=((1.03)^2)^x=(1.0609)^x\]

OpenStudy (anonymous):

why you would want to do it is anyone’s guess, but now the exponent is \(x\) instead of \(2x\)

OpenStudy (anonymous):

wait I still don't get it